OpenStudy (anonymous):

A man bought 42 stamps, some 13¢ and some 18¢. How many of each kind did he buy if the cost was $6.66?

OpenStudy (anonymous):

Let's call the $0.13 stamps T and the $0.18 stamps E (for thirteen and eighteen). T*.13 + E * .18 = 6.66 and T + E = 42 You can solve the second equation for one of the variables and then substitute; T = 42- E Then the first equation becomes (42-E)*.13 + E * .18 = 6.66

OpenStudy (anonymous):

so i got 21 of each kind?

OpenStudy (anonymous):

21 * .13 = 2.73 21 * .18 = 3.78 3.78+2.73 = 6.51...So I think we're a little off.

OpenStudy (anonymous):

5.46 - E*.13 + E*.18 = 6.66 Subtract 5.46 from both sides to get 0 - E*.13 + E*.18 = 1.20 A little rearranging and distribution property: E * (.18 - .13) = 1.20 E * (.05) = 1.20

OpenStudy (anonymous):

That should be right. I switched the variables in my head. Did that make sense? What did you get?

OpenStudy (anonymous):

Systems of Equations: x + y = 42 .13x + .18 y = 6.66 1. Solve each for y: y = 42 -x y = (6.66 - .13x)/.18 2. Set both equations equal to each other: y = y 42 - x = (6.66 - .13x)/.18 .18(42 - x) = 6.66 - .13x 7.56 - .18x = 6.66 - .13x .90 = .05x x = 18 y = 24